OpenStudy (anonymous):

to convert 0.9 to a fraction, Lauren wrote n=0.9 as her first step. For her second step, what should she multiply both sides of that equation by? the 9's have a line over them btw

OpenStudy (anonymous):

I Lost connection, sorry

OpenStudy (mathmale):

That line over the 9's indicates a repeating decimal fraction, e. g., 0.99999.... Are you positive that that is part of the "convert 0.9 to a fraction" problem? I'd suggest you write \[\frac{ 0.9 }{ 1 }\] and multiply both numerator and denominator by 10. Describe what happens to numerator and denominator. Is the result of this multiplicatino an acceptable answer, or will more work have to be done to obtain such an answer?

OpenStudy (anonymous):

Whaaa? im totally confused

OpenStudy (mathmale):

\[\frac{ 0.9 }{ 1 }*\frac{ 10 }{ 10 }=?\]

OpenStudy (anonymous):

Ummm 9/10 ?

OpenStudy (mathmale):

Again, describe what happens to 0.9 and 1 when you multiply each by 10. Is the result an acceptable fraction (with NO decimal points)? Think about this. Then reply. The goal of this problem was to transform 0.9 into a proper fraction (with NO decimal points). Good luck. I need to get off the 'Net now.

OpenStudy (anonymous):

I still dont get it...